OpenStudy (anonymous):

find the second derivative of 10sinx^2 + 20x^2cosx^2 and 7cosx^2 - 14x^2sinx^2

OpenStudy (anonymous):

once, twice, three times a derivative. do it twice, first time get \[20 \cos(x) (\sin(x)+2 x \cos(x)-2x^2\sin^2(x))\] then you have to do it again. your math teacher must hate you

OpenStudy (agreene):

Yeah, both of those are annoying to do once...much less doing a f"(x)